# Galactic transient sources with the Cherenkov Telescope Array

### v1

@@ 1.5 Magnetars
Magnetars are isolated neutron stars in which the main energy source is the magnetic field (e.g. Mereghetti et al. 2015; Kaspi & Beloborodov 2017, for reviews). They are observed as pulsed X-ray sources, with typical spin periods of a few seconds and strong spin-down rates, and/or through the detection of short bursts and flares in the hard X-ray/soft gamma-ray range. This led to their historical subdivision in the Anomalous X-ray Pulsars and Soft Gamma-ray Repeater classes (Mereghetti, 2008), but it is now clear that these are just two different manifestations of the same underlying object: a strongly magnetized neutron star powered by magnetic energy, as proposed by Paczynski (1992) and Duncan & Thompson (1992).
@@
About 30 magnetars are known so far. With the exception of two sources in the Magellanic Clouds, all of them lie in the Galactic Plane. The majority of the magnetars are transient X-ray sources that have been discovered when they became active, with an increase of their X-ray luminosity (from a quiescent level of erg s-1 up to erg s-1), accompanied by the emission of luminous and rapid bursts. This means that the total Galactic population of magnetars is larger than the currently observed sample, and more sources of this class will be known at the time of CTA observations. Furthermore, magnetar-like behavior has recently been observed in some sources originally presumed to be of a different kind, such as rotation-powered (radio) pulsars (Gavriil et al., 2008; Göğüş et al., 2016), and even in the gamma-ray binary LS I +61 303 (Torres et al., 2012; Weng et al., 2022).
@@
For what concerns the persistent emission, magnetars have not been detected above few hundred keV (Abdo et al., 2010b; Aleksić et al., 2013). Their X-ray emission typically comprises a soft thermal component that dominates in the 1-10 keV range and a hard power-law component that is believed to originate from multiple resonant scattering in the magnetosphere. The ULs in the MeV range (Li et al., 2017) indicate a turn-off of this component implying that their detectability is below the CTA capabilities, unless a different spectral component is present at higher energies. On the other hand, magnetar bursts and flares (in particular the so called Giant Flares) are potentially very interesting targets for CTA, with the only disadvantage of their unpredictable time of occurrence. Giant flares are extremely energetic and bright events, reaching isotropic peak luminosities as high as a few 1047 erg s-1 for a fraction of a second. However, they occur very rarely: only three have been seen from local magnetars in 40 years. The high luminosity of their short ( 1 s) initial peaks implies that they can be detected, with properties resembling those of short GRBs, up to distances of tens of Mpc by current hard X-ray instruments. Indeed, a few candidate extragalactic giant flares have been identified (Mazets et al., 2008; Frederiks et al., 2007; Svinkin et al., 2021; Roberts et al., 2021). Of particular interest regarding CTA’s perspective to detect giant flares is the case of the flare located in the Sculptor galaxy (NGC 253, at 3.5 Mpc) for which Fermi-LAT observation led to the detection of two high-energy photons with energies of 1.3 GeV and 1.7 GeV, likely produced via synchrotron mechanism (Fermi-LAT Collaboration et al., 2021). However, no emission from a magnetar has been yet detected at TeV energies (Abdalla et al., 2021; López-Oramas et al., 2021). For further discussion, see Section 4.6.
1.6 Other expected variable gamma-ray sources at TeV energies
In addition to the systems quoted above, it can be foreseen that CTA detects other variable gamma-ray sources that nowadays are a mere theoretical prediction. These include, for instance, runaway massive stars expelled from their formation site (del Valle & Romero, 2014). The bow shocks created by their powerful stellar winds have been proposed as sites of particle acceleration responsible for non-thermal phenomena, so far mostly observed at radio wavelengths. Significant gamma-ray variability is predicted on times scales of year as the stars move across density inhomogeneities of their surrounding medium affecting the bow shock properties.
Another kind of variable gamma-ray source based on theoretical expectations are low-mass young stellar objects such as T-Tauri stars (del Valle et al., 2011). Here, particle acceleration would be related to fast reconnection events in the stellar magnetosphere. Similarly, massive young stellar objects (MYSOs) could also produce gamma-ray emission associated with fast outflows from accretion or supersonic ejecta (see e.g. Araudo et al., 2021 and references therein). Recently, the MYSO S255 NIRS 3 has been suggested as the counterpart of a HE source (de Oña Wilhelmi et al., 2023). Perspectives for detection would be maximized at least for the closest star forming regions (see discussion by the dedicated KSP on star-forming regions, Cherenkov Telescope Array Consortium et al., 2019).
Supergiant Fast X-ray Transients (SFXTs; see Sidoli 2017 for a recent review) are high-mass X-ray binaries (HMXBs) composed of an O or B type supergiant star and a compact object; normally a neutron star (NS) is assumed, since some of them are X-ray pulsars, although formally a black hole cannot be ruled out. Their X-ray outbursts are rare (the duty cycle is less than 5 %, see Sidoli & Paizis 2018). A large dynamic range of 2 to 6 orders of magnitude (Romano et al., 2015) can be reached between quiescence and the outburst peak. Their outbursts, that can last up to a few days, are made of short (2 ks), bright X-ray flares, typically reaching 1036-1037 erg s-1. The possibility that SFXTs are MeV/TeV emitters was discussed by Walter (2007) and Sguera et al. (2009). In particular, the latter noted the spatial proximity (positional association within the error circles at gamma-ray energies, typically 1∘) of some SFXTs with AGILE and EGRET transient unidentified sources. To date, there are four unidentified gamma-ray transient sources (3EG J11225946, AGL J17343310, 3EG J18370423, AGL J20223622/VER J2019368) which are spatially correlated with members (or candidates) of the SFXT class: IGR J112155952 (Sguera, 2009), IGR J173543255 (Sguera, 2013), AX J1841.00536 (Sguera et al., 2009), and IGR J201883647 (Sguera, 2009). However, it is worth mentioning that no SFXT flare has ever been simultaneously caught at both X-rays and MeV/TeV energies, implying that, besides the spatial overlap with known SFXTs, there is no other evidence supporting a physical association, to date.
Once CTA is operational some of these objects, and other unexpected ones, will likely enrich the list of known Galactic transient emitters.
2 Sensitivity of CTA to transient detection in the Galactic plane
@@
CTA will have unprecedented sensitivity over a broad energy range and will devote a large amount of time to sources in the Galactic plane, both with a dedicated Galactic Plane Survey (GPS) (for details on the pointing strategy and expected results see CTA Consortium, 2023) and with pointed observations on specific targets. These are ideal capabilities for the discovery of new Galactic transients at TeV energies.
@@
It is then important to characterize the sensitivity of CTA in the Galactic plane. The differential sensitivity of CTA for detecting a new source is defined as the minimal flux of a source, multiplied by the mean energy squared within the given energy interval, such that the source is detectable at the significance level. It is defined within a given energy range, and for a given observation interval or exposure. We assume a test point source power-law spectral model,
@@
where we set the pivot energy, , and the spectral index, , using typical values for Galactic VHE sources. The prefactor, Pf is varied as part of the sensitivity calculation, in order to find the minimal flux value for a detection. In order to be compatible with previous analyses (e.g. Fioretti et al., 2019), we also require that the source emits at least 10 gamma-ray photons. In addition, we validate that this number of events is larger than of the corresponding contribution from backgrounds and foregrounds (cosmic rays; electrons; and other gamma-ray sources that can be coincident).
@@
We explore the performance of CTA in the Galactic plane region for various short observation intervals. For illustration, the sensitivity of the Southern array is shown in Fig. 1, considering different putative source locations. In this example, we estimated the performance for short observation intervals of 10 minutes within the energy range 100–200 GeV, exploring the detection potential of new sources in the low-energy range of CTA. We use a publicly-available Galactic sky model, based on observations of known gamma-ray sources and interstellar emission from cosmic-ray interactions in the Milky Way (CTA Consortium, 2023).888Galactic sky model available at https://zenodo.org/records/10008527 We simulate our putative transients on top of the emission derived from this sky model, such that the latter constitutes an additional background to the search.
@@
The presence of possible source variability is assessed in Section 4.1.1. Other topics such as the study of Galactic Centre sources and interstellar emission through observations of the Galactic Centre region and GPS and the prospects for the CTA and its scientific results are covered in other KSPs (see Cherenkov Telescope Array Consortium et al., 2019; CTA Consortium, 2023)
We conclude that the performance of CTA is consistent with the corresponding nominal extragalactic sensitivity (in the absence of significant foregrounds).
@@ 3 Detectability of transients of unknown origin
Apart from the transient sources of clearly identifiable type, others of unknown nature could also be serendipitously observed e.g. during a scan of the GPS. To assess the detection capabilities of CTA for such sources we simulate the populations of 100 generic transients. We consider the relatively short observation time of 1 hour (compatible with the strategy defined in CTA Consortium, 2023) during which it would be possible to detect the source and make a decision about further observations.
@@
We used the model of Yusifov & Küçük (2004) for the radial distribution of sources. The resulting spatial distribution is shown in Fig. 3. For simplicity, we did not take into account the visibility constraints, assuming that all sources are visible to either array at the time of observation.
@@
Four simulated populations are summarized in Table 1. They include different spectral shapes and parameters. For each population we employed the 0.5 h IRFs for both CTA-N and CTA-S, and also tested the Alpha configuration in the case of population 4. Both IRF sets contain three zenith angle observation options at 20∘, 40∘ and 60∘; and they also account for the azimuth dependence coming from the geomagnetic field pointing direction: North, South or an average over the azimuth direction.
For each source, we simulate the photon events list for 1 hour both for the CTA-N and CTA-S sites with a 5.0∘ ROI centered at a source, without any other sources within it, accounting only for the IRF background as seen in Fig. 1 and Fig. 2. The energy ranges used for both configurations at each site are collected in Table 2. The energy dispersion effect has been also taken into account (according to the IRFs). We then performed an unbinned maximum likelihood fitting. The test statistic (TS) equal or higher than 25 is used as criterion for a source detection.
@@
CTA-N | 0.03-200 | 0.04-200 | 0.11-200 | |||
CTA-S | 0.03-200 | 0.04-200 | 0.11-200 | |||
CTA-N (Alpha) | 0.03-200 | 0.04-200 |
CTA-S (Alpha) | 0.04-200 | 0.06-200 | 0.18-200 |
The results for all populations and observation configurations are presented in Fig. 4. We found that in the case of having a power-law spectra and fluxes in the range of ph cm-2 s-1 TeV-1, of sources for 20∘ and 40∘ zenith angles and of sources for 60∘ zenith angle will be detected, while for the population with fluxes ph cm-2 s-1 TeV-1, CTA will not detect any source during 1 hour of observation. In most cases CTA-S performs marginally better than CTA-N, with the larger difference for the North magnetic field configuration. Adding a curvature to the spectrum does not affect the detection rate in a statistically significant way. Using the Alpha Configuration (which corresponds to a first construction phase) slightly decreases the detectability, which is expected due to the reduced number of telescopes and the lack of LSTs in the CTA-S Alpha configuration (Maier et al., 2023).
4 Source detection with CTA
Galactic transients that exhibit MeV-GeV emission are specially interesting to be studied with CTA, since it is already known that non-thermal mechanisms leading to gamma-ray production are at work. We aim at understanding whether these sources of interest can also emit VHE radiation, which can be produced by the same HE mechanisms and be detected as a spectral extension, or be created by an additional component at TeV energies.
@@ 4.1 High-mass microquasars
The microquasars of the Cygnus region, Cyg X-3, Cyg X-1 and the system SS 433 are the only microquasars that have been detected in the HE regime, hence they can be considered as potential targets for the CTA observatory. After the discovery of gamma-ray emission from SS 433 above 20 TeV by HAWC (Abeysekara et al., 2018a), the CTA observations of these microquasars will be crucial to shed light on the physical mechanism responsible for the VHE emission in this type of binary systems, by investigating the limits of extreme particle acceleration in the jet.
The importance of observing this subclass of binary systems with CTA has been previously discussed in Paredes et al. (2013). In particular, a detailed study on a possible detection of a TeV flare from Cyg X-1 was presented in that paper, showing conclusions similar to our findings (see Sect. 4.1.3). In this section, we show simulations on the first microquasar detected in the VHE regime, SS 433, and estimate the detectability both of transient and persistent emission from Cyg X-3 and Cyg X-1. Even if the detection of persistent emission is not the scope of this paper, we perform this additional exercise to complement the expectations to detect microquasars with CTA. For the case of the microquasars in the Cygnus region, we carried out several CTA observation simulations by using the latest prod5-v0.1
IRFs to check if these two systems could already be detected in the first years of operation of CTA. Almost all the simulations have been carried out in the lowest range of energies for the CTA observatory, where the bulk of the emission from these binary systems is expected. For each set of observations, besides the emission from the microquasars, we simulated the main field sources of the Cygnus region: 2HWC J2006+341 (Araya & HAWC Collaboration, 2019), VER J2016+371, VER J2019+368 (Aliu et al., 2014b), Gamma Cygni SNR (Ackermann et al., 2017; Abeysekara et al., 2018b), TeV J2032+4130 (emission model as detected by MAGIC before the periastron passage of November 2017, Abeysekara et al., 2018c). This approach also applies to the case of the LMXB V404 Cyg located in the same region (see subsection 4.2.1).

### v2

@@ 1.5 Magnetars
Magnetars are isolated neutron stars in which the main energy source is the magnetic field (e.g. Mereghetti et al. 2015; Kaspi & Beloborodov 2017, for reviews). They are observed as pulsed X-ray sources, with typical spin periods of a few seconds and strong spin-down rates (typically 10-12–10-10 s s-1) (Harding et al., 1999), and/or through the detection of short bursts and flares in the hard X-ray/soft gamma-ray range. This led to their historical subdivision in the Anomalous X-ray Pulsars and Soft Gamma-ray Repeater classes (Mereghetti, 2008), but it is now clear that these are just two different manifestations of the same underlying object: a strongly magnetized neutron star powered by magnetic energy, as proposed by Paczynski (1992) and Duncan & Thompson (1992).
@@
About 30 magnetars are known so far. With the exception of two sources in the Magellanic Clouds, all of them lie in the Galactic Plane. The majority of the magnetars are transient X-ray sources that have been discovered when they became active, with an increase of their X-ray luminosity (from a quiescent level of erg s-1 up to erg s-1), accompanied by the emission of luminous and rapid bursts. This means that the total Galactic population of magnetars is larger than the currently observed sample, and more sources of this class will be known at the time of CTAO observations. Furthermore, magnetar-like behavior has recently been observed in some sources originally presumed to be of a different kind, such as rotation-powered (radio) pulsars (Gavriil et al., 2008; Göğüş et al., 2016), and even in the gamma-ray binary LS I +61 303 (Torres et al., 2012; Weng et al., 2022).
@@
For what concerns the persistent emission, magnetars have not been detected above few hundred keV (Abdo et al., 2010b; Aleksić et al., 2013). Their X-ray emission typically comprises a soft thermal component that dominates in the 1-10 keV range and a hard power-law component that is believed to originate from multiple resonant scattering in the magnetosphere. The upper limits (ULs) in the MeV range (Li et al., 2017) indicate a turn-off of this component implying that their detectability is below the CTAO capabilities, unless a different spectral component is present at higher energies. On the other hand, magnetar bursts and flares (in particular the so called Giant Flares) are potentially very interesting targets for CTAO, with the only disadvantage of their unprediCTAOble time of occurrence. Giant flares are extremely energetic and bright events, reaching isotropic peak luminosities as high as a few 1047 erg s-1 for a fraction of a second. However, they occur very rarely: only three have been seen from local magnetars in 40 years. The high luminosity of their short ( 1 s) initial peaks implies that they can be detected, with properties resembling those of short GRBs, up to distances of tens of Mpc by current hard X-ray instruments. Indeed, a few candidate extragalactic giant flares have been identified (Mazets et al., 2008; Frederiks et al., 2007; Svinkin et al., 2021; Roberts et al., 2021). Of particular interest regarding CTAO’s perspective to detect giant flares is the case of the flare located in the Sculptor galaxy (NGC 253, at 3.5 Mpc) for which Fermi-LAT observation led to the detection of two high-energy photons with energies of 1.3 GeV and 1.7 GeV, likely produced via synchrotron mechanism (Fermi-LAT Collaboration et al., 2021). However, no emission from a magnetar has been yet detected at TeV energies (Abdalla et al., 2021; López-Oramas et al., 2021). For further discussion, see Section 4.6.
2 Sensitivity of CTAO to transient detection in the Galactic plane
@@
CTAO will have unprecedented sensitivity over a broad energy range and will devote a large amount of time to sources in the Galactic plane, both with a dedicated Galactic Plane Survey (GPS) (for details on the pointing strategy and expected results see CTA Consortium, 2023) and with pointed observations on specific targets. These are ideal capabilities for the discovery of new Galactic transients at TeV energies.
@@
It is then important to characterize the sensitivity of CTAO in the Galactic plane. The differential sensitivity of CTAO for detecting a new source is defined as the minimal flux of a source, multiplied by the mean energy squared within the given energy interval, such that the source is detectable at the significance level. It is defined within a given energy range, and for a given observation interval or exposure. We assume a test point source power-law spectral model,
@@
where we set the pivot energy, , and the spectral index, , using typical values for Galactic VHE sources. The prefactor, Pf is varied as part of the sensitivity calculation, in order to find the minimal flux value for a detection. In order to be compatible with previous analyses (e.g. Fioretti et al., 2019), we also require that the source emits at least 10 gamma-ray photons. In addition, we validate that this number of events is larger than of the corresponding contribution from backgrounds (cosmic rays; electrons) and foregrounds (other coincident gamma-ray sources).
@@
We explore the performance of CTAO in the Galactic plane region for various short observation intervals. For illustration, the sensitivity of the Southern array is shown in Fig. 1, considering different putative source locations. In this example, we estimated the performance for short observation intervals of 10 minutes within the energy range 100–200 GeV, exploring the detection potential of new sources in the low-energy range of CTAO. We use a publicly-available Galactic sky model, based on observations of known gamma-ray sources and interstellar emission from cosmic-ray interactions in the Milky Way (CTA Consortium, 2023).888Galactic sky model available at https://zenodo.org/records/10008527 We simulate our putative transients on top of the emission derived from this sky model, such that the latter constitutes an additional background to the search.
@@
As a test, the presence of possible source variability is assessed in Section 4.1.1. Other topics such as the study of Galactic Centre sources and interstellar emission through observations of the Galactic Centre region and GPS and the prospects for the CTAO and its scientific results are covered in other KSPs (see Cherenkov Telescope Array Consortium et al., 2019; CTA Consortium, 2023)
We conclude that the performance of CTAO in the Galactic plane is consistent with the corresponding nominal extragalactic sensitivity. (That is, the sensitivity in the absence of significant emission from other gamma-ray sources.)
@@ 3 Detectability of transients of unknown origin
Apart from the transient sources of clearly identifiable type, others of unknown nature could also be serendipitously observed e.g. during a scan of the GPS. The detailed study of serendipity and corresponding observational strategies for CTAO will be addressed in a dedicated separate publication. However, to assess capabilities of CTAO for the detection of Galactic transient sources of unknown origin, a population of generic sources can be used. A full study of such populations requires considering various models of sources, population sizes and observational setups, and therefore will be presented separately. Here, we illustrate the methodology with a specific, simplified example. We simulate the populations of 100 generic transients. We consider the relatively short observation time of 1 hour (compatible with the strategy defined in CTA Consortium, 2023) during which it would be possible to detect the source and make a decision about further observations.
@@
We used the model of Yusifov & Küçük (2004) for the radial distribution of sources. For simplicity, we did not take into account the visibility constraints, assuming that all sources are visible to either array at the time of observation.
@@
Four simulated populations are summarized in Table 1. They include different spectral shapes and parameters. For the log-parabola model we assume the range of curvature [-0.25, 0.25]. Note that populations 1 and 4 are two different populations sharing only the spatial distribution of the sources. For each population we employed the 0.5 h IRFs for both CTAO-N and CTAO-S, and also tested the Alpha configuration in the case of population 4. Both IRF sets contain three zenith angle observation options at 20∘, 40∘ and 60∘; and they also account for the azimuth dependence coming from the geomagnetic field pointing direction: North, South or an average over the azimuth direction.
For each source, we simulate the photon events list for 1 hour both for the CTAO-N and CTAO-S sites with a 5.0∘ ROI centered at a source, without any other sources within it, accounting only for the IRF background as seen in Fig. 1 and Fig. 2. The energy ranges used for both configurations at each site are collected in Table 2. The energy dispersion effect has been also taken into account (according to the IRFs). We then performed an unbinned maximum likelihood fitting. The test statistic (TS) equal or higher than 25 is used as criterion for a source detection. The TS for different values of prefactor and spectral index are shown in Figs. 4 and 5 respectively.
@@
CTAO-N | 0.03-200 | 0.04-200 | 0.11-200 | |||
CTAO-S | 0.03-200 | 0.04-200 | 0.11-200 | |||
CTAO-N (Alpha) | 0.03-200 | 0.04-200 |
CTAO-S (Alpha) | 0.04-200 | 0.06-200 | 0.18-200 |
The results for all populations and observation configurations are presented in Fig. 3-6. Since the simulated lightcurves have different timescales and to have an idea at which lightcurve stage the observation takes place and how it affects the detectability of a source, we present in Fig. 6 the test statistic versus the value of lightcurve averaged over the observation time for each source in the simulated populations. We found that in the case of having a power-law spectra and fluxes in the range of ph cm-2 s-1 TeV-1, of sources for 20∘ and 40∘ zenith angles and of sources for 60∘ zenith angle will be detected, while for the population with fluxes ph cm-2 s-1 TeV-1, CTAO will not detect any source during 1 hour of observation. In most cases CTAO-S performs marginally better than CTAO-N, with the larger difference for the North magnetic field configuration. Adding a curvature to the spectrum does not affect the detection rate in a statistically significant way. Using the Alpha Configuration (which corresponds to a first construction phase) slightly decreases the detectability, which is expected due to the reduced number of telescopes and the lack of LSTs in the CTAO-S Alpha configuration (Maier et al., 2023).
In the case of ROIs not centered in a source, we present in Fig. 7 the dependence of the detection probability on the source distance (in degrees) from the ROI center, with a clear decrease in detectability for the offsets .
Finally, to roughly estimate how source visibility affects the presented observation probabilities, we compute, as an example, the number of detected sources for 1-h long observations starting at 2:00 UTC on three different nights taking into account the visibility constraints at each observatory site. For sources with zenith angle ranges of , and , we employed the IRFs corresponding to zenith angles of , and respectively (observations were not considered for zenith angles exceeding than ). Each source was evaluated using the IRF appropriate for its azimuth. The results are presented in Table 3. We see that while the detection probabilities without these constraints are high, there is a significant reduction in detection probabilities when visibility constraints are taken into account.
Population | IRF | 2025-05-22 | 2025-08-22 | 2025-11-22 | |||||||||||||||
zenith | CTAO-N | CTAO-S | CTAO-N | CTAO-S | CTAO-N | CTAO-S | |||||||||||||
angle | N | A | S | N | A | S | N | A | S | N | A | S | N | A | S | N | A | S | |
1 | 20∘ | 0 | 3 | 1 | 0 | 0 | 0 | 0 | 3 | 0 | 11 | 24 | 2 | 0 | 0 | 0 | 0 | 0 | 0 |
40∘ | 0 | 17 | 10 | 0 | 14 | 13 | 0 | 13 | 0 | 15 | 4 | 7 | 0 | 1 | 0 | 0 | 0 | 0 | |
60∘ | 0 | 4 | 11 | 0 | 13 | 1 | 0 | 6 | 0 | 3 | 1 | 2 | 0 | 1 | 0 | 0 | 0 | 0 | |
3 | 20∘ | 0 | 4 | 1 | 0 | 0 | 0 | 1 | 0 | 0 | 9 | 25 | 1 | 0 | 0 | 0 | 0 | 0 | 0 |
40∘ | 0 | 15 | 8 | 0 | 14 | 16 | 0 | 11 | 0 | 11 | 4 | 11 | 1 | 1 | 0 | 0 | 0 | 0 | |
60∘ | 0 | 2 | 12 | 0 | 14 | 0 | 1 | 7 | 0 | 3 | 1 | 4 | 0 | 1 | 0 | 0 | 1 | 0 | |
4 | 20∘ | 0 | 2 | 1 | 0 | 1 | 0 | 1 | 2 | 0 | 13 | 23 | 2 | 0 | 0 | 0 | 0 | 0 | 0 |
40∘ | 0 | 17 | 12 | 0 | 12 | 12 | 0 | 9 | 0 | 13 | 2 | 6 | 0 | 0 | 0 | 0 | 0 | 0 | |
60∘ | 0 | 2 | 12 | 0 | 12 | 0 | 0 | 5 | 0 | 1 | 2 | 2 | 0 | 0 | 1 | 0 | 1 | 0 |
4 Source detection with CTAO
Galactic transients that exhibit MeV-GeV emission are specially interesting to be studied with CTAO, since it is already known that non-thermal mechanisms leading to gamma-ray production are at work. We aim at understanding whether these sources of interest can also emit VHE radiation, which can be produced by the same HE mechanisms and be detected as a spectral extension, or be created by an additional component at TeV energies.
@@ 4.1 High-mass microquasars
The microquasars of the Cygnus region, Cyg X-3, Cyg X-1 and the system SS 433 are the only microquasars that have been detected in the HE regime, hence they can be considered as potential targets for the CTAO observatory. After the discovery of persistent gamma-ray emission from SS 433 above 20 TeV by HAWC (Abeysekara et al., 2018a), the latest LHAASO detections (LHAASO Collaboration, 2024) and specially the first detection by an IACT (H. E. S. S. Collaboration et al., 2024), the CTAO observations of these microquasars will be crucial to shed light on the physical mechanism responsible for the VHE emission in this type of binary systems, by investigating the limits of extreme particle acceleration in the jet.
The importance of observing this subclass of binary systems with CTAO has been previously discussed in Paredes et al. (2013). In particular, a detailed study on a possible detection of a TeV flare from Cyg X-1 was presented in that paper, showing conclusions similar to our findings (see Sect. 4.1.3). In this section, we show simulations on the first microquasar detected in the VHE regime, SS 433, and estimate the detectability both of transient and persistent emission from Cyg X-3 and Cyg X-1. Even if the detection of persistent emission is not the scope of this paper, we perform this additional exercise to complement the expectations to detect microquasars with CTAO. For the case of the microquasars in the Cygnus region, we carried out several CTAO observation simulations by using the latest prod5-v0.1
IRFs to check if these two systems could already be detected in the first years of operation of CTAO. Almost all the simulations have been carried out in the lowest range of energies for the CTAO observatory, where the bulk of the emission from these binary systems is expected. For each set of observations, besides the emission from the microquasars, we simulated the main field sources of the Cygnus region: 2HWC J2006+341 (Araya & HAWC Collaboration, 2019), VER J2016+371, VER J2019+368 (Aliu et al., 2014b), Gamma Cygni SNR (Ackermann et al., 2017; Abeysekara et al., 2018b), TeV J2032+4130 (emission model as detected by MAGIC before the periastron passage of November 2017, Abeysekara et al., 2018c). This approach also applies to the case of the LMXB V404 Cyg located in the same region (see subsection 4.2.1).